# Quantum Search with the Signless Laplacian

### v1

@@ VI Runtime Comparison
$t_{A}=\frac{\pi}{\sqrt{2}}\sqrt{\frac{N_{1}N_{2}}{k_{2}N_{1}+k_{1}N_{2}}},$
@@
where $t_{L,a}$ and $t_{L,b}$ are the runtimes for the Laplacian quantum walk search algorithm to find a marked vertex in the left or right partite set, respectively, $t_{A}$ is the runtime for the adjacency quantum walk to find a marked vertex in either partite set, and $t_{Q,a}$ and $t_{Q,b}$ are the runtimes for the signless Laplacian quantum walk search algorithm to find a marked vertex in the left or right partite set, respectively, When $N_{1}>N_{2}$, these runtimes are asymptotically
$t_{L,a}\approx\frac{\pi}{2}\sqrt{\frac{N_{1}}{k_{1}}},$
$t_{L,b}\approx\frac{\pi}{2}\sqrt{\frac{N_{1}}{k_{2}}},$
$t_{A}\approx\frac{\pi}{\sqrt{2}}\sqrt{\frac{N_{2}}{k_{2}}},$
$t_{Q,a}\approx\frac{\pi}{2}\frac{N_{1}}{\sqrt{k_{1}N_{2}}},$
$t_{Q,b}\approx\frac{\pi}{2}\sqrt{\frac{N_{2}}{k_{2}}}.$
Ranking the asymptotic runtimes that depend on $k_{1}$ from fastest to slowest,
$t_{L,a}<t_{L,Q}.$
and ranking those that depend on $k_{2}$,
$t_{Q,b}<t_{A}<t_{L,b}.$
Note [16] compared the runtimes of the Laplacian and adjacency quantum walks, but here we see that the signless Laplacian quantum walk is faster than the adjacency quantum walk, so we now compare $t_{L,a}$ and $t_{Q,b}$ to determine which is faster. We see that $t_{L,a}$ is faster when $k_{1}>k_{2}N_{1}^{2}/(N_{2}N)\approx k_{2}N_{1}/N_{2}$, $t_{Q,b}$ is faster when the inequality is flipped, and they are the same when equality is obtained. This is summarized in the first set of rows in Table 3.
@@
$N_{1}>N_{2}$ $k_{1}<k_{2}\frac{N_{1}^{2}}{N_{2}(N_{1}+N_{2})}$ $t_{Q,b}$ Fastest
$k_{1}>k_{2}\frac{N_{1}^{2}}{N_{2}(N_{1}+N_{2})}$ $t_{L,a}$ Fastest
$k_{1}=k_{2}\frac{N_{1}^{2}}{N_{2}(N_{1}+N_{2})}$ $t_{L,a}$ $t_{Q,b}$ Same
$N_{1}<N_{2}$ $k_{2}<k_{1}\frac{N_{2}^{2}}{N_{1}(N_{1}+N_{2})}$ $t_{Q,a}$ Fastest
$k_{2}>k_{1}\frac{N_{2}^{2}}{N_{1}(N_{1}+N_{2})}$ $t_{L,b}$ Fastest
$k_{2}=k_{1}\frac{N_{2}^{2}}{N_{1}(N_{1}+N_{2})}$ $t_{L,b}$ $t_{Q,a}$ Same
By swapping $a\leftrightarrow b$, $k_{1}\leftrightarrow k_{2}$, and $N_{1}\leftrightarrow N_{2}$, the runtime comparison is obtained when $N_{2}>N_{1}$, and the results are summarized in the last set of rows of Table 3.

### v2

@@ VI Runtime Comparison
$t_{A}=\frac{\pi}{\sqrt{2}}\sqrt{\frac{N_{1}N_{2}}{k_{2}N_{1}+k_{1}N_{2}}},$ (19)
@@
where $t_{L,a}$ and $t_{L,b}$ are the runtimes for the Laplacian quantum walk search algorithm to find a marked vertex in the left or right partite set, respectively, $t_{A}$ is the runtime for the adjacency quantum walk to find a marked vertex in either partite set, and $t_{Q,a}$ and $t_{Q,b}$ are the runtimes for the signless Laplacian quantum walk search algorithm to find a marked vertex in the left or right partite set, respectively.
Figure 9: Runtimes for search on a complete bipartite graph of $N_{1}=1024$ and $N_{2}=256$ vertices with $k_{1}$ varying and $k_{2}=5$ marked vertices.
@@
$N_{1}\gg N_{2}$:
$k_{1}<\frac{k_{2}N_{1}(N_{1}-N_{2})}{N_{2}N}$ $t_{Q,b}$ Fastest
$\frac{k_{2}N_{1}(N_{1}-N_{2})}{N_{2}N}<k_{1}<\frac{k_{2}N_{1}N}{N_{2}(N_{1}-N_{2})}$ $t_{A}$ Fastest
$k_{1}>\frac{k_{2}N_{1}N}{N_{2}(N_{1}-N_{2})}$ $t_{L,a}$ Fastest
$N_{1}\ll N_{2}$:
$k_{2}<\frac{k_{1}N_{2}(N_{2}-N_{1})}{N_{1}N}$ $t_{Q,a}$ Fastest
$\frac{k_{1}N_{2}(N_{2}-N_{1})}{N_{1}N}<k_{2}<\frac{k_{1}N_{2}N}{N_{1}(N_{2}-N_{1})}$ $t_{A}$ Fastest
$k_{2}>\frac{k_{1}N_{2}N}{N_{1}(N_{2}-N_{1})}$ $t_{L,b}$ Fastest
Let us compare these runtimes when $N_{1}\gg N_{2}$. First, the runtimes are plotted in Fig. 9 with varying values of $k_{1}$ when searching a complete bipartite graph where $N_{1}\gg N_{2}$. We observe that the dotted green curve ($t_{A}$) is always faster than the dashed red ($t_{L,b}$) and dot-dashed blue ($t_{Q,a}$) curves. To prove the first of these, note that $t_{A}$ is maximized when $k_{1}\to 0$, at which it takes a value of
$t_{A,\text{max}}=\frac{\pi}{\sqrt{2}}\sqrt{\frac{N_{2}}{k_{2}}},$
and this is always smaller than $t_{L,b}$ since $N_{2}<N/2$. To prove the second relationship,
$t_{Q,a}$ $=\frac{\pi}{2}\sqrt{\frac{N_{1}(N_{1}+N_{2})}{k_{1}N_{2}}}>\frac{\pi}{2}\sqrt{\frac{N_{1}(N_{1}+N_{2})}{k_{2}N_{1}+k_{1}N_{2}}}$
$>\frac{\pi}{2}\sqrt{\frac{N_{1}(N_{2}+N_{2})}{k_{2}N_{1}+k_{1}N_{2}}}=\frac{\pi}{\sqrt{2}}\sqrt{\frac{N_{1}N_{2}}{k_{2}N_{1}+k_{1}N_{2}}}$
$=t_{A},$
and so $t_{A}<t_{Q,a}$. Since $t_{A}$ is faster than both $t_{L,b}$ and $t_{Q,a}$, we can ignore the latter two in determining which algorithm is the fastest, and just compare $t_{A}$ with the remaining runtimes $t_{L,a}$ and $t_{Q,b}$. Numerically, in Fig. 9, we observe that the dot-dot-dashed orange curve ($t_{Q,b}$) is fastest for small values of $k_{1}$, the dot-dashed green curve ($t_{A}$) is fastest for intermediate values of $k_{1}$, and the solid black curve ($t_{L,a}$) is fastest for large values of $k_{1}$. Analytically, we prove this by comparing the respective runtimes from (19), yielding that $t_{Q,b}$ is faster than $t_{A}$ when $k_{1}<k_{2}N_{1}(N_{1}-N_{2})/(N_{2}N)$, and $t_{A}$ is faster than $t_{L,a}$ when $k_{1}<k_{2}N_{1}N/[N_{2}(N_{1}-N_{2})]$. As a check, for the graph in Fig. 9, these transitions occur when $k_{1}=5\cdot 1024(1024-256)/(256\cdot 1280)=12$ and $k_{1}=5\cdot 1024\cdot 1280/(256(1024-256))=100/3\approx 33.33$. These results are summarized in the first set of rows in Table 3.
By swapping $a\leftrightarrow b$, $k_{1}\leftrightarrow k_{2}$, and $N_{1}\leftrightarrow N_{2}$, the runtime comparison is obtained when $N_{2}\gg N_{1}$, and the results are summarized in the last set of rows of Table 3.